# Concurrent Composition for Differentially Private Continual Mechanisms

### v3

@@ Abstract.
Many intended uses of differential privacy involve a continual mechanism that is set up to run continuously over a long period of time, making more statistical releases as either queries come in or the dataset is updated. In this paper, we give the first general treatment of privacy against adaptive adversaries for mechanisms that support dataset updates and a variety of queries, all arbitrarily interleaved. It also models a very general notion of neighboring, that includes both event-level and user-level privacy. We prove several concurrent composition theorems for continual mechanisms, which ensure privacy even when an adversary can interleave its queries and dataset updates to the different composed mechanisms. Previous concurrent composition theorems for differential privacy were only for the case when the dataset is static, with no adaptive updates. We also give the first interactive and continual generalizations of the “parallel composition theorem” for noninteractive differential privacy. Specifically, we show that the analogue of the noninteractive parallel composition theorem holds if either there are no adaptive dataset updates or each of the composed mechanisms satisfies pure differential privacy, but it fails to hold for composing approximately differentially private mechanisms with dataset updates. Thus, we prove a tight new composition theorem for this case. In addition, we prove concurrent filter compositions theorems for the scenarios in which the privacy parameters are adaptively chosen.
We then formalize a set of general conditions on a continual mechanism $\mathcal{M}$ that runs multiple continual sub-mechanisms such that the privacy guarantees of $\mathcal{M}$ follow directly using the above concurrent composition theorems on the sub-mechanisms, without further privacy loss. This enables us to give a simpler and more modular privacy analysis of a recent continual histogram mechanism of Henzinger, Sricharan, and Steiner. In the case of approximate DP, ours is the first proof that shows that its privacy holds against adaptive adversaries.
11footnotetext: Institute of Science and Technology, Klosterneuburg, Austria (email: monika.henzinger@ista.ac.at).22footnotetext: Institute of Science and Technology, Klosterneuburg, Austria (email: roodabeh.safavi@ista.ac.at).33footnotetext: Harvard University (email: salil_vadhan@harvard.edu).
@@ 1. Introduction
Differential privacy [6] is the now-standard framework for protecting privacy when performing statistical analysis or machine learning on sensitive datasets about individuals. In addition to having a rich and rapidly growing scientific literature, differential privacy has also seen large-scale adoption by technology companies and government agencies.
The familiar formulation of differential privacy is for a noninteractive mechanism $\mathcal{M}:\mathcal{X}\rightarrow\mathcal{Y}$, which takes a dataset $x\in\mathcal{X}$ and produces a statistical release (e.g. a collection of statistics or the parameters of a machine learning model) $y\in\mathcal{Y}$. Differential privacy requires that every two datasets $x$ and $x^{\prime}$ that “differ on one individual’s data,” the output distributions $\mathcal{M}(x)$ and $\mathcal{M}(x^{\prime})$ are “close” to each other. Formally, we have a neighboring relation $\sim$ on $\mathcal{X}$, which specifies which datasets “differ on one individual’s data,” or more generally what information requires privacy protection. We say that $\mathcal{M}$ is $(\epsilon,\delta)$-DP with respect to $\sim$ if for all $x\sim x^{\prime}$, we have:
@@ 1.1. Differential Privacy Over Time
Many intended uses of differential privacy involve a system that is set up to run continuously over a long period of time, making more statistical releases as either queries come in or the dataset is updated, for example with people leaving and entering. (The latter case, with dataset updates, is known as differential privacy under continual observation, and was introduced in the concurrent works of Dwork, Naor, Pitassi, and Rothblum [7] and Chan, Shi, and Song [2].) These settings introduce a new attack surface, where an adversary may be adaptive, issuing queries or influencing dataset updates based on the responses it receives.
@@
When $\delta>0$, there are mechanisms that are differentially private against oblivious adversaries but not against adaptive ones. [^1] A simple artificial example is the following: the mechanism’s first release can include a uniformly random number $r$ from $\{1,2,\ldots,\lceil 1/\delta\rceil\}$, and then if the adversary’s next query is $r$, the mechanism publishes the entire dataset in the clear. One natural example is the “advanced composition theorem” for differentially privacy [9], which is known to fail if the adversary can select the privacy-loss parameters $(\epsilon_{i},\delta_{i})$ of the composed mechanisms adaptively [21].
@@
Until recently, adaptivity has been treated in an ad hoc manner in the differential privacy literature. For example, the aforementioned advanced composition theorem of [9] was formalized in a way that allows an adversary to adaptively choose the mechanisms and pairs of adjacent datasets, just not the privacy-loss parameters, and new composition theorems that allow for adaptive choices of privacy-loss parameters (known as privacy filters and privacy odometers) were given by [21, 25]. Other early examples where adaptive queries were analyzed include the Sparse Vector Technique (SVT) [8, 12, 18] and the Private Multiplicative Weights (PMW) algorithm [12].
@@
A few years ago, Vadhan and Wang [23] initiated a systematic study of differentially private mechanisms $\mathcal{M}$ that support adaptive queries. Specifically, they formalized the notion of an interactive mechanism $\mathcal{M}$, which begins with a dataset as input and maintains a potentially secret state as it answers queries that come in over time. The use of a secret state, as utilized in both SVT and PMW, allows for correlating the answers to queries over time, and often allows for the privacy-loss budget to degrade much more slowly than composing noninteractive mechanisms would allow. They defined $\mathcal{M}$ to be an $(\epsilon,\delta)$-DP interactive mechanism if for every pair of adjacent datasets $x\sim x^{\prime}$ and every adaptive adversary $\mathcal{A}$, the view of $\mathcal{A}$ when interacting with $\mathcal{M}(x)$ is $(\epsilon,\delta)$-indistinguishable (in the sense of Inequality (1)) from its view when interacting with $\mathcal{M}(x^{\prime})$, where the view of $\mathcal{A}$ consists of the random coin tosses of $\mathcal{A}$ and the transcript of the interaction.
@@
Vadhan and Wang [23] asked whether our existing composition theorems for differential privacy, where the mechanisms being composed are noninteractive mechanisms, extend to the concurrent composition of interactive mechanisms. In concurrent composition of interactive mechanisms, an adaptive adversary can interleave its queries to the different interactive mechanisms being composed. This is a realistic attack scenario when multiple interactive differentially private mechanisms are deployed in practice, and is also a useful primitive for analyzing differentially private algorithms that use several interactive mechanisms as building blocks.
The challenge in analyzing concurrent composition comes from the fact that the adversary can base its queries to $\mathcal{M}_{1}$ on answers given by $\mathcal{M}_{2}$ and vice-versa. The privacy guarantees on an interactive mechanism $\mathcal{M}$ only promise privacy against adversaries that do not get additional information about the sensitive dataset in the middle of its interaction with $\mathcal{M}$. Reasoning about the effect of such interleaved interactions is particularly tricky when the mechanisms $\mathcal{M}_{1}$ and/or $\mathcal{M}_{2}$ are not pure DP and can maintain secret state (like in the Sparse Vector Technique). Thus, it took several years before we had concurrent composition theorems that matched our composition theorems for noninteractive mechanisms.
@@
Note that the extension of composition theorems for noninteractive mechanism to concurrent composition theorems for interactive mechanisms is non-trivial: Let us call an interactive mechanism $M$ $M$-adaptive if it is differentially private against an adversary that was only interactive with $M$. Consider an adversary that first interacts adaptively with $M_{1}$, then $M_{2}$ and then with $M_{1}$ again. Note that for the second sequence of interactions with $M_{1}$ the adversary takes into account its prior interaction with $M_{1}$ as well as with $M_{2}$. Thus, the adversary has more information about the data than if it had only interacted with $M_{1}$ before and the fact that $M_{1}$ is $M_{1}$-adaptive does not mean that it is adaptive against this interleaved composition of $M_{1}$ and $M_{2}$.
Vadhan and Wang showed that the basic composition theorem for pure differential privacy [6] does extend to concurrent composition, but for approximate differential privacy, they obtained a much weaker bound for the concurrent case. A tight bound for concurrent composition of approximate differential privacy was given by Lyu [19]. A series of works gave concurrent composition theorems for other flavors of differential privacy and composition, such as Rényi DP [19], $f$-DP [24], and adaptive choices of privacy parameters [11].
@@
Haney, Shoemate, Tian, Vadhan, Vyrros, Wang, and Xu [11] study the concurrent extension of composition theorems for non-interactive mechanisms where the adversary instantiates arbitrary mechanisms with adaptive privacy parameters. This setting is referred to as filter composition, where a filter is a function $\mathrm{Filt}$ that maps the privacy parameters $(\epsilon_{i},\delta_{i})$ of the created mechanisms to an element of $\{\top,\bot\}$. Suppose an adversary adaptively creates noninteractive mechanisms such that $g$ applied to the privacy parameters of the created mechanisms equals $\top$ at every point in time. The $\mathrm{Filt}$-filter composition of noninteractive mechanisms is said to satisfy $(\epsilon,\delta)$-DP if the view of every such adversary is $(\epsilon,\delta)$-indistinguishable when all created mechanisms are run on dataset $x$ or its neighboring dataset $x^{\prime}$. Haney et al. extend this definition to the concurrent $\mathrm{Filt}$-filter composition of interactive mechanisms, where the adversary may not only create new mechanisms but also interact with existing interactive ones. They prove that if the $\mathrm{Filt}$-filter composition of non-interactive mechanisms is $(\epsilon,\delta)$-DP, then the concurrent $\mathrm{Filt}$-filter composition of interactive mechanisms is also $(\epsilon,\delta)$-DP.
@@
All of the above works are for the case where the queries to the mechanisms are adaptive, but the datasets are static, given as input to the mechanisms at the start of the interaction. [^2] One can extend the concurrent composition theorems to the case where, similarly to the advanced composition theorem [9], an adversary can adaptively select a pair of adjacent datasets $(x_{i},x_{i}^{\prime})$ for the $i$’th mechanism $\mathcal{M}_{i}$, but once the interactive mechanism $\mathcal{M}_{i}$ is started, its dataset is fixed and no further updates are allowed. Thus, whenever $(x_{i+1},x_{i+1}^{\prime})$ are given, a new interactive mechanism $\mathcal{M}_{i+1}$ has to be started that does not know the secret state of $\mathcal{M}_{i}$. However, in some real-world applications, both queries and dataset updates may be issued adaptively, i.e., the dynamic updates to the dataset might be chosen adaptively. For instance, Ghazi et al. [10] study an advertising problem in which the server receives both private data and queries over time. They design algorithms that accept dataset updates while applying non-interactive mechanisms—such as the discrete Laplace mechanism—to answer queries and analyze the privacy guarantees for their algorithm using existing results on privacy filters for composing non-interactive mechanisms with adaptive privacy-loss parameters. This independent work [^3] The first version of this work was available on arXiv prior to [10]. shows that if a mechanism receiving both queries and dataset updates produces a differentially private output at each time step $i$ (with adaptive parameters $\epsilon_{i},\delta_{i}$ at step $i$) such that $\sum_{i}\epsilon_{i}\leq\epsilon$ and $\sum_{i}\delta_{i}\leq\delta$, then the overall mechanism is $(\epsilon,\delta)$-DP. This guarantee can be seen as a special case of our filter composition theorem, where the composition rule is simple summation and each composed mechanism is restricted to a non-interactive mechanism corresponding to a single step. While some continual mechanisms do indeed employ independent non-interactive mechanisms at each step, many important examples—such as the Sparse Vector Technique (SVT)—generate outputs that are correlated across time. For these mechanisms, privacy must be analyzed over the entire output sequence rather than on a step-by-step basis to get the desired guarantees. (If the $\epsilon$-DP SVT mechanism were analyzed on a step-by-step basis, then the privacy guarantee would degrade to $\Theta(q\cdot\epsilon)$-DP, where $q$ is the number of queries issued, defeating the entire point of SVT, which is to have a privacy guarantee and noise scale that is independent of the number of queries.) As a consequence, in the concurrent composition of continual mechanisms (or even interactive mechanisms), one cannot simply reduce the analysis to the composition of per-step non-interactive mechanisms. Indeed, this is why concurrent composition theorems for interactive mechanisms took several years to establish [23, 19, 24, 11]. Thus, the results of Ghazi et al. [10] do not apply to the more general setting of (filter) concurrent composition of continual mechanisms. [^4] It is Ghazi et al.’s definition of Individual DP (IDP, their Definition 4.1) and their composition theorem that is framed in terms of IDP that restricts to step-by-step privacy analyses corresponding to composition of noninteractive mechanisms. Their definition of $(\epsilon,\delta)$-DP interactive mechanisms (their Definition 3.6) is somewhat more general than their definition of IDP, but it still does not allow for exploiting correlated randomness across queries to reduce privacy loss, as in SVT. (This is because their Definition 3.4 requires the next state of the mechanism to be a function of only the dataset and the query, and not of the randomness used to generate the answer.
@@
Jain, Raskhodnikova, Sivakumar, and Smith [14] recently gave a formalization of adaptive adversaries under continual observation (where there are dataset updates but the queries are fixed) for the special case of event-level privacy, where two streams of updates are considered adjacent if they differ on only one update. The lack of a more general formalism and toolkit for reasoning about adaptivity in dataset updates has led to complicated and ad hoc proofs of privacy of continual mechanisms (see e.g. [13]). Rectifying this situation is what motivated our work.
@@
Comparison with algorithms and data structures. We note that the different flavors of mechanisms discussed above are differential privacy analogues of familiar variants of randomized algorithms and data structures. A non-interactive mechanism is simply a static (or batch) algorithm. An interactive mechanism corresponds to static data structure. Our focus in this paper is on continual mechanisms, which allow both queries and updates, and these can be viewed as dynamic data structures. Adaptivity is a common concern for the correctness properties of randomized data structures; here we are concerned instead with its effect on privacy properties.
1.2. Our contributions
@@
(A) We give a formalism that captures privacy against adaptive adversaries for continual mechanisms, which support both queries and dataset updates. This generalizes the previously studied cases of interactive mechanisms (which have static datasets) and continual observation (which answer fixed queries). See Figure 1 for an illustration. We note that specific mechanisms, like variants of SVT, have been proposed to handle both dataset updates and queries [15], but without offering a definition of privacy against adversaries that can adaptively select both the updates and queries.

### v4

@@ Abstract.
Many intended uses of differential privacy involve a continual mechanism that is set up to run continuously over a long period of time, making more statistical releases as either queries come in or the dataset is updated. In this paper, we give the first general treatment of privacy against adaptive adversaries for mechanisms that support dataset updates and a variety of queries, all arbitrarily interleaved. It also models a very general notion of neighboring, that includes both event-level and user-level privacy. We prove several concurrent composition theorems for continual mechanisms, which ensure privacy even when an adversary can interleave its queries and dataset updates to the different composed mechanisms. Previous concurrent composition theorems for differential privacy were only for the case when the dataset is static, with no adaptive updates. We also give the first interactive and continual generalizations of the “parallel composition theorem” for noninteractive differential privacy. Specifically, we show that the analogue of the noninteractive parallel composition theorem holds if either there are no adaptive dataset updates or each of the composed mechanisms satisfies pure differential privacy, but it fails to hold for composing approximately differentially private mechanisms with dataset updates. Thus, we prove a tight new composition theorem for this case. In addition, we prove concurrent filter compositions theorems for the scenarios in which the privacy parameters are adaptively chosen. We extend these results to other measures of differential privacy, including Rényi DP and $f$-DP.
We then formalize a set of general conditions on a continual mechanism $\mathcal{M}$ that runs multiple continual sub-mechanisms such that the privacy guarantees of $\mathcal{M}$ follow directly using the above concurrent composition theorems on the sub-mechanisms, without further privacy loss. This enables us to give a simpler and modular privacy analysis of a recent continual histogram mechanism of Henzinger, Sricharan, and Steiner. In the case of approximate DP, ours is the first proof that shows that its privacy holds against adaptive adversaries. We also provide a framework that simplifies the analysis of local differential privacy when the protocol includes multi-round server-user interactions. Using this result, we simplify the privacy analysis of the core decomposition protocol of Dhulipala, Henzinger, Li, Liu, Sricharan, and Zhu (dhulipala2025near, ).
copyright: none
@@ 1. Introduction
Differential privacy (dwork2006calibrating, ) is the now-standard framework for protecting privacy when performing statistical analysis or machine learning on sensitive datasets about individuals. In addition to having a rich and rapidly growing scientific literature, differential privacy has also seen large-scale adoption by technology companies and government agencies.
The familiar formulation of differential privacy is for a noninteractive mechanism $\mathcal{M}:\mathcal{X}\rightarrow\mathcal{Y}$, which takes a dataset $x\in\mathcal{X}$ and produces a statistical release (e.g. a collection of statistics or the parameters of a machine learning model) $y\in\mathcal{Y}$. Differential privacy requires that for every two datasets $x$ and $x^{\prime}$ that “differ on one individual’s data,” the output distributions $\mathcal{M}(x)$ and $\mathcal{M}(x^{\prime})$ are “close” to each other. Formally, we have a neighboring relation $\sim$ on $\mathcal{X}$, which specifies which datasets “differ on one individual’s data,” or more generally what information requires privacy protection. We say that $\mathcal{M}$ is $(\epsilon,\delta)$-DP with respect to $\sim$ if for all $x\sim x^{\prime}$, we have:
@@
Differential Privacy Over Time. Many intended uses of differential privacy involve a system that is set up to run continuously over a long period of time, making more statistical releases as either queries come in or the dataset is updated, for example with people leaving and entering. (The latter case, with dataset updates, is known as differential privacy under continual observation, and was introduced in the concurrent works of Dwork, Naor, Pitassi, and Rothblum (Dwork2010, ) and Chan, Shi, and Song (DBLP:journals/tissec/ChanSS11, ) – see (DBLP:conf/esa/HenzingerS25, ) for a survey.) These settings introduce a new attack surface, where an adversary may be adaptive, issuing queries or influencing dataset updates based on the responses it receives.
@@
When $\delta>0$, there are mechanisms that are differentially private against oblivious adversaries but not against adaptive ones. [^1] A simple artificial example is the following: the mechanism’s first release can include a uniformly random number $r$ from $\{1,2,\ldots,\lceil 1/\delta\rceil\}$, and then if the adversary’s next query is $r$, the mechanism publishes the entire dataset in the clear. One natural example is the “advanced composition theorem” for differential privacy (dwork2010boosting, ), which is known to fail if the adversary can select the privacy-loss parameters $(\epsilon_{i},\delta_{i})$ of the composed mechanisms adaptively (rogers2016privacy, ).
@@
Until recently, adaptivity has been treated in an ad hoc manner in the differential privacy literature. For example, the aforementioned advanced composition theorem of (dwork2010boosting, ) was formalized in a way that allows an adversary to adaptively choose the mechanisms and pairs of adjacent datasets, just not the privacy-loss parameters, and new composition theorems that allow for adaptive choices of privacy-loss parameters (known as privacy filters and privacy odometers) were given by (rogers2016privacy, ; whitehouse2023fully, ). Other early examples where adaptive queries were analyzed include the Sparse Vector Technique (SVT) (dwork2009complexity, ; HardtRo10, ; lyu2016understanding, ) and the Private Multiplicative Weights (PMW) algorithm (HardtRo10, ).
@@
A few years ago, Vadhan and Wang (vadhan2021concurrent, ) initiated a systematic study of differentially private mechanisms $\mathcal{M}$ that support adaptive queries. Specifically, they formalized the notion of an interactive mechanism $\mathcal{M}$, which begins with a dataset as input and maintains a potentially secret state as it answers queries that come in over time. The use of a secret state, as utilized in both SVT and PMW, allows for correlating the answers to queries over time, and often allows for the privacy-loss budget to degrade much more slowly than composing noninteractive mechanisms would allow. They defined $\mathcal{M}$ to be an $(\epsilon,\delta)$-DP interactive mechanism if for every pair of adjacent datasets $x\sim x^{\prime}$ and every adaptive adversary $\mathcal{A}$, the view of $\mathcal{A}$ when interacting with $\mathcal{M}(x)$ is $(\epsilon,\delta)$-indistinguishable (in the sense of Inequality (1)) from its view when interacting with $\mathcal{M}(x^{\prime})$, where the view of $\mathcal{A}$ consists of the random coin tosses of $\mathcal{A}$ and the transcript of the interaction.
@@
Vadhan and Wang (vadhan2021concurrent, ) asked whether our existing composition theorems for differential privacy, where the mechanisms being composed are noninteractive mechanisms, extend to the concurrent composition of interactive mechanisms. In concurrent composition of interactive mechanisms, an adaptive adversary can interleave its queries to the different interactive mechanisms being composed. This is a realistic attack scenario when multiple interactive differentially private mechanisms are deployed in practice, and is also a useful primitive for analyzing differentially private algorithms that use several interactive mechanisms as building blocks. Vadhan and Wang showed that the basic composition theorem for pure differential privacy (dwork2006calibrating, ) extends to concurrent composition. This result was later generalized by Lyu (lyu2022composition, ) to all composition theorems for approximate differential privacy. Subsequent works gave concurrent composition theorems for other flavors of differential privacy, such as Rényi DP (lyu2022composition, ) and $f$-DP (vadhan2022concurrent, ).
@@
Note that the extension of composition theorems for noninteractive mechanisms to concurrent composition theorems for interactive mechanisms is non-trivial: Let us call an interactive mechanism $M$ $M$-adaptive private if it is differentially private against an adversary that was only interactive with $M$. Consider an adversary that first interacts adaptively with $M_{1}$, then $M_{2}$ and then with $M_{1}$ again. Note that for the second sequence of interactions with $M_{1}$ the adversary takes into account its prior interaction with $M_{1}$ as well as with $M_{2}$. Thus, the adversary has more information about the data than if it had only interacted with $M_{1}$ before and the fact that $M_{1}$ is $M_{1}$-adaptive private does not mean that it is private against this interleaved composition of $M_{1}$ and $M_{2}$.
@@
Haney, Shoemate, Tian, Vadhan, Vyrros, Wang, and Xu (haney2023concurrent, ) study the concurrent extension of composition theorems for noninteractive mechanisms where the adversary instantiates arbitrary mechanisms with adaptive privacy parameters. This setting is referred to as filter composition, where a filter is a function $\mathrm{Filt}$ that maps the privacy parameters $(\epsilon_{i},\delta_{i})$ of the created mechanisms to an element of $\{\top,\bot\}$. Suppose an adversary adaptively creates noninteractive mechanisms such that $\mathrm{Filt}$ applied to the privacy parameters of the created mechanisms equals $\top$ at every point in time. The $\mathrm{Filt}$-filter composition of noninteractive mechanisms is said to satisfy $(\epsilon,\delta)$-DP if the view of every such adversary is $(\epsilon,\delta)$-indistinguishable when all created mechanisms are run on dataset $x$ or its neighboring dataset $x^{\prime}$. Haney et al. extend this definition to the concurrent $\mathrm{Filt}$-filter composition of interactive mechanisms, where the adversary may not only create new mechanisms but also interact with existing interactive ones. They prove that if the $\mathrm{Filt}$-filter composition of noninteractive mechanisms is $(\epsilon,\delta)$-DP, then the concurrent $\mathrm{Filt}$-filter composition of interactive mechanisms is also $(\epsilon,\delta)$-DP.
@@
All of the above works are for the case where the queries to the mechanisms are adaptive, but the datasets are static, given as input to the mechanisms at the start of the interaction. [^2] One can extend the concurrent composition theorems to the case where, similarly to the advanced composition theorem (dwork2010boosting, ), an adversary can adaptively select a pair of adjacent datasets $(x_{i},x_{i}^{\prime})$ for the $i$’th mechanism $\mathcal{M}_{i}$, but once the interactive mechanism $\mathcal{M}_{i}$ is started, its dataset is fixed and no further updates are allowed. Thus, whenever $(x_{i+1},x_{i+1}^{\prime})$ are given, a new interactive mechanism $\mathcal{M}_{i+1}$ has to be started that does not know the secret state of $\mathcal{M}_{i}$. However, in some real-world applications, both queries and dataset updates may be issued adaptively, i.e., the dynamic updates to the dataset might be chosen adaptively. For instance, Ghazi et al. (ghazi2025privacy, ) study an advertising problem in which the server receives both private data and queries over time. They design algorithms that accept dataset updates while applying noninteractive mechanisms—such as the discrete Laplace mechanism—to answer queries and analyze the privacy guarantees for their algorithm using existing results on privacy filters for composing noninteractive mechanisms with adaptive privacy-loss parameters. This independent work [^3] The first version of this work was available on arXiv prior to (ghazi2025privacy, ). shows that if a mechanism receiving both queries and dataset updates produces a differentially private output at each time step $i$ (with adaptive parameters $\epsilon_{i},\delta_{i}$ at step $i$) such that $\sum_{i}\epsilon_{i}\leq\epsilon$ and $\sum_{i}\delta_{i}\leq\delta$, then the overall mechanism is $(\epsilon,\delta)$-DP. This guarantee can be seen as a special case of our filter composition theorem, where the composition rule is simple summation and each composed mechanism is restricted to a noninteractive mechanism corresponding to a single step. While some continual mechanisms do indeed employ independent noninteractive mechanisms at each step, many important examples—such as the Sparse Vector Technique (SVT)—generate outputs that are correlated across time. For these mechanisms, privacy must be analyzed over the entire output sequence rather than on a step-by-step basis to get the desired guarantees. (If the $\epsilon$-DP SVT mechanism were analyzed on a step-by-step basis, then the privacy guarantee would degrade to $\Theta(q\cdot\epsilon)$-DP, where $q$ is the number of queries issued, defeating the entire point of SVT, which is to have a privacy guarantee and noise scale that is independent of the number of queries.) As a consequence, in the concurrent composition of continual mechanisms (or even interactive mechanisms), one cannot simply reduce the analysis to the composition of per-step noninteractive mechanisms. Indeed, this is why concurrent composition theorems for interactive mechanisms took several years to establish (vadhan2021concurrent, ; lyu2022composition, ; vadhan2022concurrent, ; haney2023concurrent, ). Thus, the results of Ghazi et al. (ghazi2025privacy, ) do not apply to the more general setting of (filter) concurrent composition of continual mechanisms. [^4] It is Ghazi et al.’s definition of Individual DP (IDP, their Definition 4.1) and their composition theorem that is framed in terms of IDP that restricts to step-by-step privacy analyses corresponding to composition of noninteractive mechanisms. Their definition of $(\epsilon,\delta)$-DP interactive mechanisms (their Definition 3.6) is somewhat more general than their definition of IDP, but it still does not allow for exploiting correlated randomness across queries to reduce privacy loss, as in SVT. (This is because their Definition 3.4 requires the next state of the mechanism to be a function of only the dataset and the query, and not of the randomness used to generate the answer.)
@@
Jain, Raskhodnikova, Sivakumar, and Smith (pricedpco, ) recently gave a formalization of adaptive adversaries under continual observation (where there are dataset updates but the queries are fixed) for the special case of event-level privacy, where two streams of updates are considered adjacent if they differ on only one update. Denisov, McMahan, Rush, Smith and Thakurta (denisov2022improved, ) extended this definition to arbitrary neighbor relations on update sequences. The lack of a general toolkit for reasoning about adaptivity in dataset updates has led to complicated and ad hoc proofs of privacy of continual mechanisms (see e.g. (henzingerSS2023, )). Rectifying this situation is what motivated our work.
@@
Comparison with algorithms and data structures. We note that the different flavors of mechanisms discussed above are differential privacy analogues of familiar variants of randomized algorithms and data structures. A noninteractive mechanism is simply a static (or batch) algorithm. An interactive mechanism corresponds to a static data structure. Our focus in this paper is on continual mechanisms, which allow both queries and updates, and these can be viewed as dynamic data structures. Adaptivity is a common concern for the correctness properties of randomized data structures; here we are concerned instead with its effect on privacy properties.
@@
Our contributions (A) We give a formalism that captures privacy against adaptive adversaries for continual mechanisms, which support both queries and dataset updates (Definition 3.7). This generalizes the previously studied cases of interactive mechanisms (which have static datasets) and continual observation (where mechanisms answer fixed queries). See Figure 4 in Appendix A for an illustration. We note that specific mechanisms, like variants of SVT, have been proposed to handle both dataset updates and queries (JainSmWa24, ), but without offering a definition of privacy against adversaries that can adaptively select both the updates and queries.
(B) We prove concurrent composition theorems for continual mechanisms, where an adaptive adversary may interleave queries and updates across mechanisms based on the entire history of observed outputs. As discussed above for the concurrent composition of interactive mechanisms, the concurrent setting involves additional adaptivity that is not supported by composition theorems for noninteractive mechanisms. [^5] Existing results on the concurrent composition of interactive mechanisms do not apply to the concurrent composition of continual mechanisms, since in our model the adversary adaptively selects both (a) the mechanisms receiving neighboring inputs (including their privacy parameters and neighboring definition) and (b) the neighboring update sequences they receive over time, which is not the case for interactive mechanisms. We show tight privacy bounds for the following useful forms of concurrent composition of continual mechanisms. Further explanations are provided in Appendix B.
  1. (1)
    We show that when the privacy-loss parameters are fixed, the standard composition theorems for noninteractive approximate differential privacy extend to the concurrent composition of continual mechanisms (Theorem 4.12). This result also holds for Rényi differential privacy (Corollary 13.4) and for $f$-DP [^6] The $f$-DP results in this paper assume finite communication: for every created continual mechanism, there exists a constant $C$ bounding the sizes of its query and answer spaces and the number of queries it can answer before halting. This assumption was also made for interactive mechanisms in (vadhan2022concurrent, ). (Corollary 14.7).
  1. (2)
    We extend the classical “parallel composition theorem” for noninteractive mechanisms to the concurrent parallel composition of an unbounded [^7] Throughout this paper, we use the term “unbounded number” to refer to a quantity of unknown (potentially arbitrary) size. number of adaptively chosen interactive mechanisms with an adaptively chosen pair of neighboring datasets (Corollary 5.15). We show similar extensions for Rényi differential privacy (Corollary 13.7) and for $f$-DP (Corollary 14.9).
  2. (3)
    We extend the above concurrent parallel composition of interactive mechanisms to continual mechanisms when the composed mechanisms satisfy pure differential privacy (Corollary 5.5). For approximate DP, we construct a counterexample demonstrating that the concurrent parallel composition of an unbounded number of continual mechanisms with adaptively chosen dataset updates is not private (Theorem 5.2). Since $(\epsilon,\delta)$-DP is a special case of $f$-DP, this lower bound also applies to $f$-DP. We further construct an analogous counterexample for the case where the composed mechanisms satisfy Rényi DP instead of approximate DP (Theorem 13.5).
  3. (4)
    By imposing an upper bound on $\prod_{i}(1-\delta_{i})$, where $\delta_{i}$ is the approximate privacy parameter of the $i$-th created mechanism, we derive a new tight composition theorem for concurrent parallel composition of continual mechanisms (Theorem 5.4).
  4. (5)
    As an immediate consequence, our new theorem extends the classical parallel composition theorem for noninteractive mechanisms to the concurrent composition of an unbounded number of adaptively chosen continual mechanisms satisfying pure differential privacy (Corollary 5.5).
  5. (6)
    We show that, for every filter $\mathrm{Filt}$, the concurrent $\mathrm{Filt}$-filter composition of continual mechanisms enjoys the same privacy guarantees as the $\mathrm{Filt}$-filter composition of noninteractive mechanisms (Theorem 12.4). This result is also proved for Rényi DP (Corollary 13.8) and $f$-DP (Corollary 14.10).
A summary of the positive results for the aforementioned types of concurrent composition of $(\epsilon,\delta)$-DP interactive and continual mechanisms is provided in Table 1.
Concurrent Composition with Fixed Parameters Concurrent Parallel Composition Concurrent Filter Composition
CMs
Theorem 4.12 for $(\epsilon,\delta)$-DP
Corollary 5.5 for $\epsilon$-DP
Theorem 5.4 for $(\epsilon,\delta)$-DP
Theorem 12.4 for $(\epsilon,\delta)$-DP
IMs
(vadhan2021concurrent, ) for $\epsilon$-DP
(vadhan2022concurrent, ; lyu2022composition, ) for $(\epsilon,\delta)$-DP
Corollary 5.15 for $(\epsilon,\delta)$-DP
(haney2023concurrent, ) for $(\epsilon,\delta)$-DP
Table 1. Concurrent composition results for $(\epsilon,\delta)$-DP interactive mechanisms (IMs) and continual mechanisms (CMs).